0.3x-y=0
y=2+0.25x
solve by emlimination

Answer by ewatrrr(24785)


Hi
0.3x-y=0, y=2+0.25x
solving by elimination
0.3x -y = 0
.25x -y = -2
0.3x -y = 0
-.25x +y = 2
.05x = 2
x = 40 and y = 12 (2 + 10)
CHECKING our Answer***
.3*40 - 12 = 12 -12 = 0